If (xn) and (yn) are sequences in the same normed space X, show that xn→x and yn→y implies xn+yn→x+y as well as αxn→αx, where α is any scalar.


1
Expert's answer
2021-09-28T08:45:46-0400

The set of all sequences in the normed space X is a vector space under the natural component-wise definitions of vector addition and scalar multiplication:

"\{x_1,x_2,...\}+\{y_1,y_2,...\}=\{x_1+y_1,x_2+y_2,...\}" and

"\alpha\{x_1,x_2,...\}=\{\alpha x_1,\alpha x_2,...\}"

So, if  xn→x and yn→y, then:

xn+yn→x+y and αxn→αx, where α is any scalar.